OpenStudy (anonymous):

factor 16(m +3)^2-25

OpenStudy (anonymous):

difference of 2 squares...if you cant see that try substitution. let x=m+3 16x^2-25=(4x-5)(4x+5) substitute m+3 back in (4(m+3)-5)(4(m+3)+5)=(4m+7)(4m+17)